OpenStudy (anonymous):

Let f(x) = 2x - 6 Solve f-1(x) when x = 2 im confused will give medal

OpenStudy (anonymous):

just plug in 2 for every x and then solve

OpenStudy (anonymous):

ok so i got f(2)=2(2)-6 f(2)=4-6

OpenStudy (anonymous):

how would i solve after that